Circular No. 8079

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003ax AND 2003ay Further to IAUC 8077, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of more apparent supernovae on NEAT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003ax Feb. 9.3 8^{h}08^{m}35.^{s}53 + 7^{o}46′42′′.2 18.7 4′′ W, 1′′ N
2003ay Feb. 19.1 4 07 26.25 +28 07 55.4 19.1 2′′ W, 7′′ N

Additional magnitudes: SN 2003ax, Jan. 26 UT, [19.5; Feb. 18, 18.4; 19, 18.6; 22, 18.8; 23, 19.0. SN 2003ay, Jan. 28, [20.0; Feb. 22, 19.1; 23, 19.2.

POSSIBLE METEORS FROM COMET C/1976 D1 P. Jenniskens, SETI Institute, Mountain View, CA, reports that cal-

culations by E. Lyytinen and himself (Icarus, in press) indicate that a one-revolution dust trail of comet C/1976 D1 (Bradfield; see also IAUC 2933) will be in the earth’s path on 2003 Mar. 1.912 ± 0.010 UT for an estimated duration of 14 min (FWHM) with a peak rate of perhaps 300/hr. The trail passes just inside the earth’s orbit at a distance of 0.00008 AU. The predicted radiant is at α = 13^{o}, δ = −64^{o} (equinox 2000.0), and the apparent meteor velocity would be 42 km/s.

COMET C/2002 Y1 (JUELS-HOLVORCEM) Ephemeris extension to IAUC 8055 (elements MPC 47728; H_{15} = 6.5):

2003 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 2 21^{h}26.^{m}60 +57^{o}30.^{′}2 0.977 1.088 67.^{o}1 57.^{o}0 7.0
7 22 14.34 +53 39.8 1.040 1.021 60.3 57.6 6.7
12 22 48.89 +49 30.2 1.111 0.957 53.7 56.8 6.4
17 23 14.38 +45 22.9 1.188 0.896 47.5 54.9 6.2
22 23 33.84 +41 25.0 1.265 0.841 41.6 51.8 5.9
27 23 49.27 +37 37.0 1.341 0.794 36.1 47.7 5.6
Apr. 1 0 01.97 +33 56.6 1.411 0.756 30.9 42.8 5.4
6 0 12.87 +30 21.3 1.474 0.729 26.4 37.6 5.3
11 0 22.61 +26 49.0 1.527 0.715 22.7 32.7 5.2

Visual m_{1} estimates: Feb. 6.68 UT, 8.4 (Y. Nagai, Yamanashi, Japan, 11×80 binoculars); 12.19, 7.9 (A. Baransky, Bucha, Ukraine, 20×60 binoculars); 20.93, 7.7 (J. J. Gonzalez, Asturias, Spain, 11×80 binoculars); 23.77, 7.4 (M. Meyer, Riedelbach, Germany, 10×50 binoculars).

2003 February 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green